OpenStudy (scarlettfarra2000):

The variables y varies inversely with x and y=8 when x=4 what is y when x=2? A) 4 B) 5 C) 16 D) 24

OpenStudy (welshfella):

Y = k x Where k is a constant to be found You can find k by plugging in y= 8 and x = 4 into the above formula

OpenStudy (welshfella):

- and solve for k

OpenStudy (welshfella):

Then you can get the required formula by replacing k by the value you just found. Nex plug in x = 2 and find the required value of y.

OpenStudy (scarlettfarra2000):

If I'm solving for k it would like something like \[y=\frac{ k }{ 4 }?\]

OpenStudy (welshfella):

Yes sorry I misread the question It Inverse variation so it's y = k / x

OpenStudy (welshfella):

So we have 8 = k / 4

OpenStudy (scarlettfarra2000):

so the answer would end up being 16 right?

OpenStudy (welshfella):

Yes K = 32 So y = 32 / 2 = 16

OpenStudy (scarlettfarra2000):

thank you for your help

OpenStudy (welshfella):

Yw